# CURVALID: Geometrically-guided Adversarial Prompt Detection

### v1

@@ B.2.7 Replacement of CNN in CurvaLID Step 1 with Transformer and RNN models
B.2.8 Performance of CurvaLID on adversarial prompts with reordered word sequences
@@
B.2.9 Performance of CurvaLID on reordered persuasive social-engineered adversarial prompts
In this section, we evaluate CurvaLID on PAIR, DAN, and Persuasive attacks, all of which are social-engineered persuasive attacks designed to preserve both semantic meaning and adversarial intent while varying structure (Chao et al., 2023; Shen et al., 2023; Zeng et al., 2024). To introduce linguistic variations, we utilized GPT-4-o to reorder the words in each sentence of the adversarial prompts while maintaining their semantic meaning. The experimental setup remains the same as described in B.2.1, with the addition of 300 DAN prompts. The experimental results, presented in Table 14, show that CurvaLID consistently achieved over 96% detection accuracy across all three attack types and maintained a 0% attack success rate on Vicuna. These findings highlight the robustness of our method, even against sophisticated and linguistically varied prompts specifically crafted to bypass defenses. This robustness underscores CurvaLID’s potential for deployment in real-world scenarios, where adversarial prompts are likely to exploit linguistic diversity to evade detection.
@@ B.2.9 Performance of CurvaLID on reordered persuasive social-engineered adversarial prompts
B.2.10 Differences in PromptLID and TextCurv between benign and adversarial prompts under linguistic reordering
@@
B.2.11 CurvaLID with separated benign training and testing data
@@
B.2.12 CurvaLID with long text length benign prompts
We conducted an additional experiment to evaluate CurvaLID’s performance on benign prompts with longer text lengths. Specifically, we calculated the median text length of benign prompts (106 characters in our experiment), removed all benign prompts with fewer than the median, and reevaluated CurvaLID’s performance as outlined in Section 5.1 (Page 7, Line 351). The results, presented in Table 17 below, show that this adjustment had minimal effect on detection accuracy. The overall accuracy was 0.990, compared to 0.992 when all benign prompts (without filtering by text length) were included. This confirms that text length has minimal impact on CurvaLID’s performance, highlighting its robustness and ability to generalize across prompts of varying lengths. Such adaptability makes CurvaLID particularly well-suited for real-world applications, where input lengths can vary significantly.
@@ B.2.12 CurvaLID with long text length benign prompts
B.2.13 CurvaLID with non-standard benign prompts
@@
B.2.14 Analysis of PromptLID and TextCurv distributions
@@ B.3.1 LID Estimation using Method of Moments
- Compute the mean of the first $k-1$ nearest distances:
@@ B.3.2 Top 10 most common nearest-neighbors for different prompts in different datasets
Dataset Top 10 most common nearest neighbors SAP ␣to, ␣and, ␣the, <.>, ␣a, ␣of, <,>, ␣<”>, ␣that, ␣your DAN <,>, <.>, <newline>, ␣the, ␣and, ␣to, ␣you, ␣will, ␣not, ␣is MathAttack ␣the, <,>, ␣of, ␣he, <.>, ␣to, ␣and, ␣she, ␣is, ␣a GCG ␣text, tto, use, ized, <}>, ␣mar, dt, ␣a, <’>, ␣Guide Orca ␣the, <,>, <.>, <newline>, ␣and, ␣a, ␣of, ␣to, ␣is, ␣in MMLU ␣the, <,>, <.>, ␣of, ␣a, ␣to, ␣that, ␣and, ␣in, ␣was AlpacaEval ␣the, ␣a, ␣to, <,>, ␣of, ␣and, ␣I, <.>, ␣for, <?> TQA ␣the, <?>, ␣a, ␣you, ␣is, ␣of, ␣that, ␣to, ␣if, ␣in
@@ B.3.4 Top 10 most common Nearest Neighbor words after removing stop words and punctuation
B.4 Performance of other SOTA defenses
We summarize the performance of various defenses across different LLMs, focusing on their ability to reduce the ASR of adversarial prompts and compare this to CurvaLID’s unified performance. CurvaLID outperforms the studied defenses and maintains consistent performance across all LLMs. While we primarily compare four key defenses in this subsection—SmoothLLM (Robey et al., 2023), IntentionAnalysis (Zhang et al., 2024a), RTT3d (Yung et al., 2024), and LAT (Sheshadri et al., 2024)—which are considered SOTA or represent some of the most recent developments, we also experimented with other defenses like Gradient Cuff (Hu et al., 2024), SELFDEFEND (Wang et al., 2024), and SafeDecoding (Xu et al., 2024a).
Given the computational and time constraints associated with replicating results and testing across multiple LLMs, we have cited the performance figures for these defenses from their respective papers. This approach ensures fairness, as different studies report varying results for these defenses when replicating them against different adversarial prompts and across different models. Therefore, we rely on the original reported performances to provide a balanced and consistent comparison. Note that since the results for these defenses are primarily based on English adversarial prompts in their respective papers, our analysis here is focused solely on English prompts.
Defenses based on input perturbation demonstrate mixed results depending on the LLM and the nature of the adversarial attack. For instance, IntentionAnalysis reduces the ASR to between 0.03% and 8.34%, but it struggles with models like Vicuna-7B and MPT-30B-Chat, where the ASR for the SAP attack can reach nearly 20%. Similarly, while SmoothLLM can reduce the ASR to nearly 0% in various LLMs, it fails against PAIR attacks, showing ASRs of 46% in Vicuna-13B and 24% in GPT-4. RTT3d, as the first defense against MathAttack, managed to mitigate 40% of MathAttack in GPT4, but it failed to reduce the ASR to under 10%. LAT achieves near-zero ASR for models like Llama2-7B-Chat and Llama3-8B-Instruct. However, its reliance on a white-box setting and its testing on models with fewer than 10 billion parameters limit its broader applicability.
In the remainder of this subsection, we present the defensive performance of the SOTA defenses. The performance figures for all defenses are directly cited from their original papers due to the computational and time constraints involved in replicating results and testing across various LLMs. This approach ensures consistency and fairness in comparison, as the results reported by different studies often vary when replicating these defenses on different adversarial prompts and models. By relying on the figures from the original sources, we aim to provide an accurate and balanced reflection of each defense’s performance.
We first compared the defenses using a common LLM, LLaMA-2-7B-Chat, as all the studied defenses employ this model. Table 22 presents the ASR comparison between CurvaLID and seven other defenses, showcasing CurvaLID’s exceptional performance in mitigating various adversarial prompts.
Table 22: ASR comparison of CurvaLID with other defenses against different adversarial prompts on LLaMA-2-7B-Chat. A dash (-) indicates that experiments were not conducted for this setting in the original paper.
@@ B.4 Performance of other SOTA defenses
LLM Defense GCG PAIR DAN AmpleGCG SAP MathAttack RandomSearch LLaMA-2-7B-Chat SmoothLLM 0.1 8 - 0 - - 0 LAT 0.007 0.025 - - - - - Gradient Cuff 0.012 0.23 - - - - - IntentionAnalysis 0 - 0.13 - 0 - - SELFDEFEND 0.09 0.21 0.242 - - - - RTT3d 0.17 0.043 - - - - - SafeDecoding 0 0.04 - - - - - CurvaLID (ours) 0 0 0 0.025 0 0 0
@@
The following are the experimental settings and performances of the seven defenses we studied. Note that it includes various LLMs, namely Vicuna (Chiang et al., 2023), Llama-2 (Touvron et al., 2023), Llama-3 (AI@Meta, 2024), GPT-3.5 (Brown, 2020), GPT-4 (Achiam et al., 2023), PaLM2 (Anil et al., 2023), Claude-1 (Anthropic, 2023a), Claude-2 (Anthropic, 2023b), ChatGLM-6B (Zeng et al., 2022), MPT-30B-Chat (Team, 2023), DeepSeek-67B-Chatand (Bi et al., 2024). It also includes various adversarial prompts, namely, GCG (Zou et al., 2023), PAIR (Chao et al., 2023), DAN (Shen et al., 2023), AmpleGCG (Liao & Sun, 2024), SAP (Deng et al., 2023a), MathAttack (Zhou et al., 2024b), RandomSearch (Andriushchenko et al., 2024), Prefill (Haizelabs, 2023), Many-Shot (Anil et al., 2024), AutoDAN (Liu et al., 2023), TAP (Mehrotra et al., 2023), Jailbroken (Wei et al., 2024), LRL (Yong et al., 2023), DrAttack (Li et al., 2024), Puzzler (Chang et al., 2024), MultiJail (Deng et al., 2023b), DeepInception (Li et al., 2023b), and Template (Yu et al., 2023).
@@
Table 23: ASR comparison of different LLMs against adversarial prompts under SmoothLLM defense. A dash (-) indicates that experiments were not conducted for this setting in the original paper.
@@
Table 24: ASR comparison of different LLMs against adversarial prompts under LAT defense.
@@
Table 25: ASR comparison of different LLMs against adversarial prompts under Gradient Cuff defense.
@@
Table 26: ASR comparison of different LLMs against adversarial prompts under IntentionAnalysis defense. A dash (-) indicates that experiments were not conducted for this setting in the original paper.
@@
Table 27: ASR comparison of different LLMs against adversarial prompts under SELFDEFEND defense.
@@
Table 28: ASR comparison of different LLMs against adversarial prompts under RTT3d defense. A dash (-) indicates that experiments were not conducted for this setting in the original paper.
@@
Table 29: ASR comparison of different LLMs against adversarial prompts under SafeDecoding defense.

### v2

@@ B.2.17 Replacement of CNN in CurvaLID Step 1 with Transformer and RNN models
B.2.18 Performance of CurvaLID on adversarial prompts with reordered word sequences
@@
B.2.19 Performance of CurvaLID on reordered persuasive social-engineered adversarial prompts
In this section, we evaluate CurvaLID on PAIR, DAN, and Persuasive attacks, all of which are social-engineered persuasive attacks designed to preserve both semantic meaning and adversarial intent while varying structure (Chao et al., 2023; Shen et al., 2023; Zeng et al., 2024). To introduce linguistic variations, we utilized GPT-4-o to reorder the words in each sentence of the adversarial prompts while maintaining their semantic meaning. The experimental setup remains the same as described in B.2.1, with the addition of 300 DAN prompts. The experimental results, presented in Table 25, show that CurvaLID consistently achieved over 96% detection accuracy across all three attack types and maintained a 0% attack success rate on Vicuna. These findings highlight the robustness of our method, even against sophisticated and linguistically varied prompts specifically crafted to bypass defences. This robustness underscores CurvaLID’s potential for deployment in real-world scenarios, where adversarial prompts are likely to exploit linguistic diversity to evade detection.
@@ B.2.19 Performance of CurvaLID on reordered persuasive social-engineered adversarial prompts
B.2.20 Differences in PromptLID and TextCurv between benign and adversarial prompts under linguistic reordering
@@
B.2.21 CurvaLID with separated benign training and testing data
@@
B.2.22 CurvaLID with long text length benign prompts
We conducted an additional experiment to evaluate CurvaLID’s performance on benign prompts with longer text lengths. Specifically, we calculated the median text length of benign prompts (106 characters in our experiment), removed all benign prompts with fewer than the median, and reevaluated CurvaLID’s performance. The results, presented in Table 28 below, show that this adjustment had minimal effect on detection accuracy. The overall accuracy was 0.990, compared to 0.992 when all benign prompts (without filtering by text length) were included. This confirms that text length has minimal impact on CurvaLID’s performance, highlighting its robustness and ability to generalize across prompts of varying lengths. Such adaptability makes CurvaLID particularly well-suited for real-world applications, where input lengths can vary significantly.
@@ B.2.22 CurvaLID with long text length benign prompts
B.2.23 CurvaLID with non-standard benign prompts
@@
B.2.24 CurvaLID on held-out datasets (out-of-distribution data) and prompt-length robustness
We performed two additional robustness evaluations. First, we ran a held-out dataset study across four benign and four adversarial datasets (200 prompts per dataset). For each run, CurvaLID was trained on seven datasets and evaluated on the remaining one. Table 30 reports per-dataset accuracy on the held-out sets, showing consistently high performance (around 0.9), indicating strong out-of-distribution generalization.
Table 30: CurvaLID detection accuracy on held-out datasets for OOD evaluation.
Dataset Orca MMLU AlpacaEval TruthfulQA GCG PAIR DAN AmpleGCG Accuracy 0.955 0.94 0.945 0.995 1.0 0.875 1.0 1.0
Second, we assessed robustness to prompt length by training on prompts shorter than 106 characters (the median benign length) and testing on longer prompts. As summarized in Table 31, the overall accuracy drops by only 0.05, suggesting minimal sensitivity to prompt length.
Table 31: CurvaLID performance under OOD evaluation on prompt length.
Model Benign Acc. Adv. Acc. Overall Acc. CurvaLID trained with shorter prompts 0.911 0.965 0.938 CurvaLID 0.984 1.000 0.992
B.2.25 Amplification of TextCurv differences through CNN activation
We investigated how TextCurv differs between benign and adversarial prompts across CNN layers in CurvaLID. As shown in Table 32, we observe that CNN activation significantly amplifies the curvature gap: the mean TextCurv of adversarial prompts is at least 30% higher than that of benign prompts in both convolutional layers. In contrast, when calculated using only the word embeddings, the difference is notably smaller—4.91 for benign prompts versus 5.42 for adversarial prompts—amounting to a 13% increase, less than half of the gap observed in the CNN layers. All results are averaged over 10 independent runs using different random seeds. The experimental settings remain consistent with those described in Appendix B.1.5.
Table 32: Mean TextCurv values of benign and adversarial prompts based on word embeddings only and across CNN layer representations in CurvaLID.
Word Embedding Embedding Only Conv Layer 1 Conv Layer 2 Benign Adv. Benign Adv. Benign Adv. RoBERTa 4.91 5.42 (+13.0%) 0.626 0.881 (+40.7%) 0.325 0.446 (+37.2%)
B.2.26 analysis of PromptLID and TextCurv distributions
@@ B.2.25 Amplification of TextCurv differences through CNN activation
B.2.27 Effectiveness of global intrinsic dimension in adversarial prompt detection.
Figure 8: Comparison of average prompt length and global intrinsic dimension (GID) across datasets.
Global intrinsic dimension (GID) is another plausible approach for the word-level representation. It can avoid aggregating the LID for each word by assessing the GID for each word within the prompt and output a single value. We use the MLE-based estimate from Tulchinskii et al. (Tulchinskii et al., 2024). However, as shown in Figure 8, GID shows no clear distinction between benign and adversarial datasets. Instead, it strongly correlates with prompt length, with Pearson and Spearman correlation coefficients of 0.92 and 0.98, respectively. Even after removing stop words and punctuation, the results were similar, with a Pearson correlation coefficient of 0.9, highlighting the limitations of GID in detecting adversarial prompts.
@@ B.3.1 LID Estimation using Method of Moments
- Compute the mean of the first k1k-1 nearest distances:
@@ B.3.2 Top 10 most common nearest-neighbors for different prompts in different datasets
Dataset Top 10 most common nearest neighbors SAP ␣to, ␣and, ␣the, <.>, ␣a, ␣of, <,>, ␣<">, ␣that, ␣your DAN <,>, <.>, <newline>, ␣the, ␣and, ␣to, ␣you, ␣will, ␣not, ␣is MathAttack ␣the, <,>, ␣of, ␣he, <.>, ␣to, ␣and, ␣she, ␣is, ␣a GCG ␣text, tto, use, ized, <}>, ␣mar, dt, ␣a, <’>, ␣Guide Orca ␣the, <,>, <.>, <newline>, ␣and, ␣a, ␣of, ␣to, ␣is, ␣in MMLU ␣the, <,>, <.>, ␣of, ␣a, ␣to, ␣that, ␣and, ␣in, ␣was AlpacaEval ␣the, ␣a, ␣to, <,>, ␣of, ␣and, ␣I, <.>, ␣for, <?> TQA ␣the, <?>, ␣a, ␣you, ␣is, ␣of, ␣that, ␣to, ␣if, ␣in
@@ B.3.4 Top 10 most common Nearest Neighbor words after removing stop words and punctuation
B.4 Performance of other SOTA defences
To ensure a comprehensive evaluation, we organize this section into two parts. In Section B.4.1, we present results from running SOTA defences on our local environment, providing a consistent and fair comparison against CurvaLID. In Section B.4.2, we summarize the reported performances of other SOTA defences cited from their respective papers, offering a broader context across different LLMs.
B.4.1 Evaluation of SOTA defences
In this subsection, we first present the results of evaluating five baseline defences—SmoothLLM (Robey et al., 2023), Self-Reminder (Xie et al., 2023), Intentionanalysis (Zhang et al., 2024a), In-Context Demonstration defence (ICD) (Wei et al., 2023), and RTT3d (Yung et al., 2024)—alongside our proposed CurvaLID. We evaluated these methods across four LLMs: Vicuna-7B-v1.1 (Chiang et al., 2023), Llama2-7B-Chat (Touvron et al., 2023), GPT-3.5 (Brown, 2020), and PaLM2 (Anil et al., 2023), and against seven different types of adversarial prompts: GCG (Zou et al., 2023), PAIR (Chao et al., 2023), DAN (Shen et al., 2023), AmpleGCG (Liao & Sun, 2024), SAP (Deng et al., 2023a), MathAttack (Zhou et al., 2024b), and RandomSearch (Andriushchenko et al., 2024). The experiment setting follows B.1.5 and the word embedding used for CurvaLID is RoBERTa. The experimental results, measured by ASR (%), are reported in Tables 36. CurvaLID consistently outperforms the baseline defences across most scenarios. We emphasize that all results are obtained through independent evaluation under a unified experimental setup.
Table 36: We compare CurvaLID with state-of-the-art defences, including SmoothLLM (Robey et al., 2023), Self-Reminder (Xie et al., 2023), Intentionanalysis (Zhang et al., 2024a), ICD (Wei et al., 2023), and RTT3d (Yung et al., 2024). The best results are boldfaced.
@@ B.4.1 Evaluation of SOTA defences
LLM defence GCG PAIR DAN AmpleGCG SAP MathAttack RandomSearch Vicuna-7B No defence 86.0 98.0 44.5 98.0 69.0 24.0 94.0 SmoothLLM 5.5 52.0 13.0 4.2 44.6 22.0 48.5 Self-Reminder 9.5 48.0 35.5 11.5 25.2 22.0 6.0 Intentionanalysis 0.0 8.5 3.3 0.3 0.23 20.0 0.0 ICD 0.2 5.2 40.4 0.9 32.8 22.0 0.2 RTT3d 0.2 0.3 22.0 3.5 33.5 20.2 2.5 CurvaLID 0.0 0.0 0.0 1.1 0.0 0.0 0.0 LLaMA2-7B No defence 12.5 19.0 2.0 81.0 9.5 11.7 90.0 SmoothLLM 0.0 11.0 0.2 0.2 1.2 11.2 0.0 Self-Reminder 0.0 8.0 0.3 0.0 0.0 11.1 0.0 Intentionanalysis 0.0 5.8 0.7 0.0 0.0 11.2 0.0 ICD 0.0 2.7 0.8 0.0 0.0 10.8 0.0 RTT3d 0.2 0.2 1.8 0.4 5.5 9.8 0.8 CurvaLID 0.0 0.0 0.0 0.0 0.0 0.0 0.0 GPT-3.5 No defence 12.0 48.0 6.33 82.0 0.9 10.5 73.0 SmoothLLM 0.0 4.9 0.3 0.7 0.0 10.9 0.0 Self-Reminder 0.0 0.9 2.5 0.3 0.1 11.2 0.0 Intentionanalysis 0.0 0.9 0.8 0.3 0.0 10.0 0.0 ICD 0.0 0.7 0.9 0.4 0.4 9.9 0.0 RTT3d 0.3 0.2 0.8 0.2 0.7 7.6 0.0 CurvaLID 0.0 0.0 0.0 0.0 0.0 0.0 0.0 PaLM2 No defence 14.9 98.0 49.7 88.9 55.1 18.9 91.9 SmoothLLM 5.5 38.7 6.7 7.2 41.2 9.8 45.3 Self-Reminder 2.3 36.7 22.3 4.7 21.4 13.3 3.7 Intentionanalysis 0.0 2.3 1.3 0.9 0.0 9.7 0.0 ICD 0.1 4.9 34.2 0.2 33.9 9.3 0.0 RTT3d 0.1 0.1 25.5 3.3 25.0 10.2 2.8 CurvaLID 0.0 0.0 0.0 0.0 0.0 0.0 0.0
We also evaluated constrained SFT(Qi et al., 2025), using the fine-tuned Gemma-2-9B model released by the authors on GitHub and HuggingFace. The experimental results are shown in Table 37. We observe that constrained SFT is effective in mitigating attacks that rely on gibberish prefixes or suffixes, such as GCG and AmpleGCG, but fails to defend against social-engineering-based attacks like PAIR, DAN, and SAP. Most importantly, CurvaLID outperforms the constrained SFT defence across all adversarial attack types.
Table 37: Comparison of defences on Gemma-2-9B, measured by ASR (%).
LLM defence GCG PAIR DAN AmpleGCG SAP MathAttack RandomSearch Gemma-2-9B No defence 90.2 23.8 80.0 81.3 44.5 22.8 94.5 Constrained SFT 22.5 28.5 83.5 29.2 78.8 22.0 90.0 CurvaLID 0.0 0.0 0.0 2.1 0.0 0.0 0.0
B.4.2 Reported performance of existing defences
We summarize the reported performances of other SOTA defences cited from their respective papers across different LLMs, focusing on their ability to reduce the ASR of adversarial prompts and compare this to CurvaLID’s unified performance. CurvaLID outperforms the studied defences and maintains consistent performance across all LLMs. While we primarily compare four key defences in this subsection—SmoothLLM (Robey et al., 2023), Intentionanalysis (Zhang et al., 2024a), RTT3d (Yung et al., 2024), and LAT (Sheshadri et al., 2024)—which are considered SOTA or represent some of the most recent developments, we also experimented with other defences like Gradient Cuff (Hu et al., 2024), SELFDEFEND (Wang et al., 2024), SafeDecoding (Xu et al., 2024a), Circuit Breakers (Zou et al., 2024), Llama Guard (Inan et al., 2023), and perplexity-based filtering (Alon & Kamfonas, 2023).
Given the computational and time constraints associated with replicating results and testing across multiple LLMs, we have cited the performance figures for these defences from their respective papers. This approach ensures fairness, as different studies report varying results for these defences when replicating them against different adversarial prompts and across different models. Therefore, we rely on the original reported performances to provide a balanced and consistent comparison. Note that since the results for these defences are primarily based on English adversarial prompts in their respective papers, our analysis here is focused solely on English prompts.
defences based on input perturbation demonstrate mixed results depending on the LLM and the nature of the adversarial attack. For instance, Intentionanalysis reduces the ASR to between 0.03% and 8.34%, but it struggles with models like Vicuna-7B and MPT-30B-Chat, where the ASR for the SAP attack can reach nearly 20%. Similarly, while SmoothLLM can reduce the ASR to nearly 0% in various LLMs, it fails against PAIR attacks, showing ASRs of 46% in Vicuna-13B and 24% in GPT-4. RTT3d, as the first defence against MathAttack, managed to mitigate 40% of MathAttack in GPT4, but it failed to reduce the ASR to under 10%. LAT achieves near-zero ASR for models like Llama2-7B-Chat and Llama3-8B-Instruct. However, its reliance on a white-box setting and its testing on models with fewer than 10 billion parameters limit its broader applicability.
In the remainder of this subsection, we present the defensive performance of the SOTA defences. The performance figures for all defences are directly cited from their original papers due to the computational and time constraints involved in replicating results and testing across various LLMs. This approach ensures consistency and fairness in comparison, as the results reported by different studies often vary when replicating these defences on different adversarial prompts and models. By relying on the figures from the original sources, we aim to provide an accurate and balanced reflection of each defence’s performance.
@@ B.4.2 Reported performance of existing defences
The following are the experimental settings and performances of the seven defences we studied. Note that it includes various LLMs, namely Vicuna (Chiang et al., 2023), Llama-2 (Touvron et al., 2023), Llama-3 (AI@Meta, 2024), GPT-3.5 (Brown, 2020), GPT-4 (Achiam et al., 2023), PaLM2 (Anil et al., 2023), Claude-1 (Anthropic, 2023a), Claude-2 (Anthropic, 2023b), ChatGLM-6B (Zeng et al., 2022), MPT-30B-Chat (Team, 2023), DeepSeek-67B-Chatand (Bi et al., 2024). It also includes various adversarial prompts, namely, GCG (Zou et al., 2023), PAIR (Chao et al., 2023), DAN (Shen et al., 2023), AmpleGCG (Liao & Sun, 2024), SAP (Deng et al., 2023a), MathAttack (Zhou et al., 2024b), RandomSearch (Andriushchenko et al., 2024), Prefill (Haizelabs, 2023), Many-Shot (Anil et al., 2024), AutoDAN (Liu et al., 2023), TAP (Mehrotra et al., 2023), Jailbroken (Wei et al., 2024), LRL (Yong et al., 2023), DrAttack (Li et al., 2024), Puzzler (Chang et al., 2024), MultiJail (Deng et al., 2023b), DeepInception (Li et al., 2023a), and Template (Yu et al., 2023).
@@
Table 38: ASR comparison of different LLMs against adversarial prompts under SmoothLLM defence. Results are directly cited from the original paper. A dash (-) indicates that experiments were not conducted for this setting in the original paper.
@@
Table 39: ASR comparison of different LLMs against adversarial prompts under LAT defence. Results are directly cited from the original paper.
@@
Table 40: ASR comparison of different LLMs against adversarial prompts under Gradient Cuff defence. Results are directly cited from the original paper.
@@
Table 41: ASR comparison of different LLMs against adversarial prompts under Intentionanalysis defence. Results are directly cited from the original paper. A dash (-) indicates that experiments were not conducted for this setting in the original paper.
@@
Table 42: ASR comparison of different LLMs against adversarial prompts under SELFDEFEND defence. Results are directly cited from the original paper.
@@
Table 43: ASR comparison of different LLMs against adversarial prompts under RTT3d defence. Results are directly cited from the original paper. A dash (-) indicates that experiments were not conducted for this setting in the original paper.
@@
Table 44: ASR comparison of different LLMs against adversarial prompts under SafeDecoding defence. Results are directly cited from the original paper.
@@
Circuit Breakers Detailed experimental settings are referred to (Zou et al., 2024). The experimental results are shown in Table 45.
Table 45: Comparison between CurvaLID and Circuit Breakers on LLaMA-3-8B. Results are reported in terms of ASR (%).
LLM defence Adversarial Prompt GCG PAIR DAN AmpleGCG SAP RandomSearch LLaMA-3-8B Circuit Breakers 2 3.33 0 2.5 0 0 CurvaLID 0 0 0 2.5 0 0
Llama Guard Detailed experimental settings are referred to (Inan et al., 2023). The experimental results are shown in Table 46.
Table 46: Comparison of LLaMA Guard 3 and CurvaLID. Results are reported in terms of ASR (%).
defence PAIR DAN SAP AutoDAN GCG AmpleGCG MMLU AlpacaEval LLaMA Guard 3 0 0 0 0 0 0 0.09 0.01 CurvaLID 0 0 0 0 0 0.025 0 0.02
Perplexity-based filtering Detailed experimental settings are referred to (Alon & Kamfonas, 2023). The experimental results are shown in Table 47.
Table 47: Comparison of Perplexity Filtering and CurvaLID. Results are reported in terms of ASR (%).
defence GCG AmpleGCG DAN AutoDAN SAP PAIR Persuasive Attack Perplexity Filtering 0 0 0.475 0.38 1 0.624 0.76 CurvaLID 0 0.015 0 0 0 0 0
Appendix C LLM Usage
This research directly concerns LLMs, and all experiments necessarily involved their usage. In addition, we used LLMs in a limited capacity to aid and polish the writing of this paper.